OpenStudy (anonymous):

The expression 4n – 1 is a rule to describe a certain sequence. What are the first three terms of the sequence? three, seven, eleven four, eight, twelve three, six, nine four, seven, eleven

OpenStudy (anonymous):

four , seven , eleven

OpenStudy (anonymous):

can you explain how you got that answer ?

OpenStudy (anonymous):

okay the expression 4n-1 , the coefficient 4 shows it is going to repeat , 4 , 8 , 12 , and so on . but there is a (-1) , that shows you are taking away one for each ,except 4 . see how option A says 3,7,11 that is not correct because there is a 3 in the beggining , do you understand?

OpenStudy (anonymous):

so basically you start with 4 , 8 ,12 , but then you minus 1 from 8 & 12 , so that makes it , four seven eleven

OpenStudy (anonymous):

oh okay. thank you for explaining !

OpenStudy (anonymous):

but , since you're here.. can you just check to see if i did this other one right ?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

okay. so this next one I picked 7+n.

OpenStudy (anonymous):

Which expression describes the rule of this sequence? 7, 14, 21, 28, 35… 7 + n 7n 7n n+ 7 + 7

OpenStudy (anonymous):

yes that is right :) are on on FLVS?

OpenStudy (anonymous):

yes :( hbu?

OpenStudy (anonymous):

lol yea -_- what grade?

OpenStudy (anonymous):

7thgr. hbu?

OpenStudy (anonymous):

9th

OpenStudy (anonymous):

aha. I hate it. it's too stressful .

OpenStudy (anonymous):

aha yeah i agree with you on that one . Did you ever submit this test?

OpenStudy (anonymous):

not yet. I have 2 more questions left.

OpenStudy (anonymous):

okay good luck :)

OpenStudy (anonymous):

Ehh. Thank you.

OpenStudy (anonymous):

dont submit yet

OpenStudy (anonymous):

im checking over my first answer i gave you.

OpenStudy (anonymous):

it is 3,7,11

OpenStudy (anonymous):

sorry so A for your 1st question

OpenStudy (anonymous):

Which expression describes the rule of this sequence? Three, six, nine, twelve, fifteen, eighteen, and so on. three plus the quantity n minus one three times n three plus the quantity three times n three plus n

OpenStudy (anonymous):

thats A?

OpenStudy (anonymous):

no The expression 4n – 1 is a rule to describe a certain sequence. What are the first three terms of the sequence? three, seven, eleven four, eight, twelve three, six, nine four, seven, eleven thats A

OpenStudy (anonymous):

you sure ?

OpenStudy (anonymous):

yees positive

OpenStudy (anonymous):

would you like me to explain how i got this ?

OpenStudy (anonymous):

Hm. It doesn't matter. If you would like to you can.

OpenStudy (anonymous):

well so you understand for next time , take 4n-1 , plug in the values 1,2,3 , so 4*1-1 =3 , 4*2-1=7 , 4*3-1=11 .

OpenStudy (anonymous):

your plugging in 1,2,3 for the value of n

OpenStudy (anonymous):

aha. okay

OpenStudy (anonymous):

thanks.

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

have you done flvs for middle school?